. A man lifts a box weighing 56N through a vertical height 1.3 m and then walks horizontally for 2.2 m at constant speed while holding the box. What is the work done by the student? *

1
Expert's answer
2022-04-11T12:06:25-0400

The work done by the student

"W=W_1+W_2\\
=mgh+0=56*1.3=73\:\rm J"